A car travels at a constant speed of 15m/s. How many miles does the car travel in 15 sec
Answer by Theo(13342)

rate * time = distance
15 miles per second times 15 seconds equals 225 miles.
15 miles per second is equivalent to 15 * 60 = 900 miles per minute which is equivalent to 54000 miles per hour.
no car travels that fast.
no plane travels that fast.
maybe a rocket travels that fast after it leaves the gravitational pull of the earth.
this indicates something wrong with your problem.
i think you might have meant the car is traveling at 15 miles per hour.
if the car travels 15 miles per hour then the car travels 15 / 60 = .25 miles per minute and travels .25 / 60 = .004167 miles per second.
since the car is traveling at .004167 miles per second, then the car is traveling .004167 * 15 = .0625 miles in 15 seconds.
.0625 miles in 15 seconds is equivalent to .0625 * 5280 = 330 feet because there are 5280 feet in a mile.